17 + 34 + 51 + ........+ 595 =
Answer by Alan3354(69443)

17 + 34 + 51 + ........+ 595 =
-------
= 17*(1 + 2 + 3 ... + 35)
-----
(0 + 1 + 2 + 3 ... + 35) is 36 integers
Make 18 pairs:
0+35
1+34
2+33
etc
Each pair is 35
35*18 = 630
----
17*630 = 10710